Dec 19, 2008:

12:16 PM Processing/CalibSite edited by mark1
11:05 AM Processing/CalibSite created by mark1
11:05 AM Processing edited by mark1

Dec 18, 2008:

5:35 PM Ticket #208 (Boresight, flight day 330/2008, Little Rissington) created by benj
Data location: …
5:03 PM Ticket #207 (Read file info from .hdr file when using -Bs in azgcorr) created by mggr
When run with external bil/bsq files, azgcorr currently requires a -B …
3:24 PM Processing/LeverArmOffset edited by benj
Updated due to previous incorrect information about direction of gamma … (diff)

Dec 17, 2008:

5:22 PM Ticket #206 (VOC_05, flight day 307/2008, Chile) created by jatt
Data location: …
3:06 PM Ticket #205 (VOC_04, flight day 305/2008, Chile) created by jatt
Data location: …
2:56 PM Sensors/LeicaLIDAR/MashUp edited by mggr
12:07 PM Ticket #203 (Support: 8/Dec/2008, Turkia Almoustafa, GB08/03) closed by mggr
fixed: Checked with her - everything is apparently fine now.
11:02 AM Ticket #46 (Investigate PARGE) closed by benj
fixed: Did some trials but between me and Daniel Schlapfer at rese being away …
10:57 AM Ticket #129 (Support: 18/Apr/2008, Luke Bateson, BGS07/02) closed by benj
fixed: Had a look at vent areas (see attached images) to see if I could see …
10:51 AM Ticket #29 (GB2007/07, flight day 102a/2007, Inverclyde) closed by benj
fixed: Check was fine, re-delivered via FTP on 09/10/08 (but forgot to …
10:49 AM Ticket #166 (Support: 13/Jun/08, John Stevenson, IPY07/02, Lidar interpolation issue) closed by benj
fixed: This was handled by Gabriel Amable, I don't think we ever heard what …
10:39 AM Ticket #194 (Azgcorr ENVI header issues (Support: Karl Hennermann, GB08/20)) closed by benj
fixed: Most recent versions of azgcorr and azexhdf seem to have this working. …
10:35 AM Ticket #199 (Unable to re-interpolate nav on old level 1 HDF files) closed by benj
fixed: Following discussions with Bill, it has been decided that the easiest …

Dec 11, 2008:

10:36 AM Ticket #201 (Support: 5/Nov/2008, Li Zhang, no code (old data)) closed by mggr
fixed: Lily is going to work with the positioning parameters then use GCPs. …

Dec 9, 2008:

5:20 PM Ticket #204 (Support: 8/Dec/2008, Abdulwasey Mohammed, GB06/07) created by mggr
Initial contact from Abdulwasey, who was having problems processing …

Dec 8, 2008:

6:46 PM Ticket #203 (Support: 8/Dec/2008, Turkia Almoustafa, GB08/03) created by mggr
Turkia having problems with azexhdf dumping core, attached file …
12:09 PM Ticket #202 (Central collection of ground data) created by mggr
Currently when ground data is taken, it is only held by the PI/team …
11:19 AM Ticket #201 (Support: 5/Nov/2008, Li Zhang, no code (old data)) created by mggr
Contacted by Dr Li (Lily) Zhang re: some CASI data from '97 and …

Dec 5, 2008:

11:17 AM Ticket #200 (Atmospheric correction investigation) created by mggr
Per Dec 2008's steering committee, we need to investigate the methods …

Dec 4, 2008:

12:07 PM Ticket #75 (Acquire 2008 radiometric calibration) closed by mggr
fixed: Detailed report on the attempt at Reports. Short version - …
10:43 AM Reports edited by benj
Added July 2008 calibration report (diff)
10:41 AM interim_calibration_report_200807.doc attached to Reports by benj
Preliminary calibration report for July 2008 calibration
Note: See TracTimeline for information about the timeline view.